Massachusetts Institute of Technology Handout 0 6.042J/18.062J: Mathematics for Computer Science September 7, 2000 Professors Nancy Lynch and Shlomi Dolev

Tutorial Signup Form

Please complete the form below and

Return this sheet to us at the end of lecture today,

so that we can complete your assignment to a tutorial section.

Last Name, First Name, Middle Name:

E-mail Address:

Each 6.042 student must attend a two-hour tutorial on Fridays, beginning tomorrow, September 8. We will assign students to tutorial sections this evening. Section assignments will be posted on the course web page (http://theory.lcs.mit.edu/classes/6.042), by some reasonable hour this evening.

To help us plan these assignments, please

• cross out those time slots that conflict with your nondiscretionary commitments (e.g., another course, team practice), and

• circle your preferred time slots; you may write numbers beneath the circles to indicate your order of preference.

9–11  10–12  11–1  1–3  2–4  3–5